OpenStudy (albert0898):

Help with these 3 questions -- please explain.

OpenStudy (albert0898):

1. \[5\sqrt{200xy^2}\] 2. \[\sqrt[3]{40a^4}\] 3. \[\sqrt[4]{48a^9b^3}\]

OpenStudy (anonymous):

Do you know how to simplify radicals?

OpenStudy (anonymous):

Are you finding the square root of the first one or the fifth root?